How to solve a depreciation problem.
A car cost $20,000 new,depreciation formula is 20,000(0.8)^x
x being time in years. How many years until car is worth half it's original value

Answer by rwm(914)

10000=20,000(0.8)^x
1=2(0.8)^x
1/2=.8^x
3.1=x